# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

### v1

@@ 3.3. Proof of Theorem 1
Under the variable change $-q=e^{iu}$, we have
@@
The diagonal terms of $\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}$ have been described in [37, Section 2]:
$\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}=\sum_{r=1}^{\infty}\left(\frac{r}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{1}{2}\frac{(-q)+1}{(-q)-1}\right)\alpha_{-r}\alpha_{r}+\ldots,$
@@
After putting together (3.4), (3.5) and (3.6) and shifting $h=g-1$, we obtain:
@@
after $-q=e^{iu}$. The connected/disconnected calculus of Proposition 10 then yields
@@ 3.4. Proof of Proposition 10
  • $\mathbf{k}=(k_{1},\ldots,k_{s})$ is a partition of $n-m$,
  • $\mathbf{g}=(g_{0},\ldots,g_{s})$ is a partition of $g+s-1$,
@@
¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{1}}^{\circ}(\pi,(1^{k_{1}}))\times_{\overline{\mathcal{M}}_{1,1}}\ldots\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}})),$
¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times\overline{\mathcal{M}}^{\circ}_{g_{1}}(\pi,(1^{k_{1}}))\times\ldots\times\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}}))\,,$
where ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ is the base change of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ along the diagonal $\Delta:\overline{\mathcal{M}}_{1,1}\to\overline{\mathcal{M}}_{1,1}\times\ldots\times\overline{\mathcal{M}}_{1,1}$.
The fixed locus of ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},n,(2,1^{n-2}))$ for the action of the torus $(\mathbb{C}^{*})^{2}$ is the disjoint union of the stacks ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$, up to an automorphism factor which we denote by $\operatorname{Aut}(\mathbf{k},\mathbf{g})$, and the restriction of the virtual class [¯$[\overline{\mathcal{M}}_{g}^{\bullet}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}$ to ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ agrees with $\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}$. By the localization formula of [18], we have
[¯$[\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\frac{\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}}{e_{(\mathbb{C}^{*})^{2}}(N^{vir})}\,,$
where the contribution of the virtual normal bundle on each factor of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ is given by
(3.8) $\begin{split}\frac{1}{\mathsf{e}_{(\mathbb{C}^{*})^{2}}(N^{vir})}&=\frac{c(\mathbb{E}^{\vee}\otimes t_{1})c(\mathbb{E}^{\vee}\otimes t_{2})}{t_{1}t_{2}}\\ &=-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-1}+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-2}-(t_{1}+t_{2})\lambda_{g_{i}-1}\lambda_{g_{i}-2}+\ldots\end{split}$
if $g_{i}>1$, and $-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{1}+1$ if $g_{i}=1$. We denote the class (3.8) by $N_{g_{i}}$, and note that the lower order terms in the Hodge classes will not contribute for dimensional reasons. After integration, we obtain
@@
where $\operatorname{ev}_{i}:\overline{\mathcal{M}}_{(m,\mathbf{k},\mathbf{g})}\to\overline{\mathcal{M}}_{1,1}$ recovers the isomorphism class of the target curve of the $i$-th map. The second equality follows from the decomposition of the diagonal of ¯$\overline{\mathcal{M}}_{1,1}$:
$[\Delta]=\sum_{j=0}^{s}[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\boxtimes 1\boxtimes[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\in\operatorname{\mathsf{CH}}^{s}(\overline{\mathcal{M}}_{1,1}\times\cdots\times\overline{\mathcal{M}}_{1,1})\,.$
@@
We now follow the geometric notation and analysis of Section 2.5. The moduli space of stable maps to the fibers of $\pi$ lies over ¯$\overline{\mathcal{M}}_{1,1}$,
@@
$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\times_{\overline{\mathcal{M}}_{1,1}}\mathcal{E}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\,.$$h$$\nu$$\pi_{M}$
@@
over ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$. Since $k_{i}>0$, the map (3.10) is injective, so we obtain an exact sequence
@@
where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$. We therefore have a factorization
@@
on ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$, where $[E]\in\overline{\mathcal{M}}_{1,1}$ is the moduli point of a fixed nonsingular elliptic curve $E$.
@@
(3.12) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})]^{vir}}\lambda_{g_{i}}\lambda_{g_{i}-1}=\frac{1}{24}\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(E,k_{i})]^{vir}}\lambda^{2}_{g_{i}-1}\,.$
@@
Therefore, the only positive contributions to the right hand side of (3.9) come from triplets $(e,\mathbf{k},\mathbf{g})$ where $\mathbf{g}=(g,1,\ldots,1)$.
@@ 3.5. Proof of Proposition 8
where the gluing identifies the section $\mathsf{p}_{1}$ of $\mathcal{E}$ with the 00-section of $\mathbb{P}(T_{\mathsf{p}_{1}}\mathcal{E}\oplus\mathbb{C})$. The universal section $\mathsf{p}_{1}$ is now the $\infty$-section of $\pi_{P}$. We will use the degeneration formula
@@ 1. $\bullet$
    [¯$[\overline{\mathcal{M}}_{g}(\pi/\mathsf{p}_{1})]^{vir}\cap\operatorname{ev}_{\overline{\mathcal{M}}_{1,1}}^{*}(\mathsf{pt})=[\overline{\mathcal{M}}_{g}(E/p_{1})]^{vir}$
@@ 2. $\bullet$
    Vertex (3.21) vanishes because of the two point conditions over ¯$\overline{\mathcal{M}}_{1,1}$.
@@ 3. $\bullet$
The vertices (3.18) and (3.20) contribution factors $1$ and $1/2$ respectively (the latter is cancelled by the multiplicity which occurs in the degeneration formula).
@@ 4.2. Proof of Theorem 4
We adapt the reconstruction strategy of [50, Section 6] to prove the result. Let $m\geq 0$, and fix classes
@@
Pulling back Getzler’s relation from ¯$\overline{\mathcal{M}}_{1,4}$ to ¯$\overline{\mathcal{M}}_{1,4+m}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),d)$ and summing over the curve degree $d$, we obtain the following equation (see [26, Equation (27)])
[…]
where the relation $\mathsf{X}\sim\mathsf{Y}$ means that the difference $\mathsf{X}-\mathsf{Y}$ is an explicit sum of arbitrary genus 00 invariants and genus $1$ invariants with at most $m+1$ insertions. Getzler’s relation can be equivalently cast in the following form:
@@
from (4.2). A linear algebraic result [50, Lemma 6.2] about solutions to the difference equation then yields
@@
Repeated use of relation (4.7) shows that $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$ can be expressed in terms of genus 00 invariants and $1$-point genus $1$-functions. Moreover, since we use the divisor equation in the argument, the resulting expressions are linear in $1$-point genus $1$ functions and their $q\frac{d}{dq}$ derivatives. By the genus 00 reconstruction result in [37], the genus 00 series can be written as rational functions in matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$. The proof of Theorem 4 is complete. ∎
@@ 3. $\bullet$
Since the eigenvectors $v_{i}$ are found by solving the equation $D\star_{\mathsf{t}}v_{i}=e_{i}v_{i}$, the components of $v_{i}$ are rational functions of $e_{i}$ with coefficients in the matrix coefficients of $D\star_{\mathsf{t}}$. Therefore, the $\Delta_{i}$ are rational functions in the eigenvalues $e_{i}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$, where the matrix coefficients of $D\star_{\mathsf{t}}$ lie.
To evaluate the term $\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})$, we need only evaluate symmetric rational functions in the derivatives of the eigenvalues $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$. By Proposition 14 of Appendix A, these expressions lie in the field of rational functions of derivatives of the symmetric functions of $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$. The outcome is an explicit calculation of
$\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
@@
$\nabla_{v_{1}}\nabla_{v_{2}}\cdots\nabla_{v_{k}}\log(\prod_{i}\Delta_{i})\,\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$

### v2

@@ 3.3. Proof of Theorem 1
Under the variable change $q=e^{iu}$, we have
@@
The diagonal terms of $\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}$ have been described in [39, Section 2]:
$\frac{1}{t_{1}+t_{2}}\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}=\sum_{r=1}^{\infty}\left(\frac{r}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{1}{2}\frac{(-q)+1}{(-q)-1}\right)\alpha_{-r}\alpha_{r}+\ldots,$
@@
After putting together (3.4), (3.5) and (3.6) and shifting h=g1h=g-1, we obtain:
@@
after $q=e^{iu}$. The connected/disconnected calculus of Proposition 10 then yields
@@ 3.4. Proof of Proposition 10
  • $\mathbf{k}=(k_{1},\ldots,k_{s})$ is a partition of nmn-m,
  • $\mathbf{g}=(g_{0},\ldots,g_{s})$ is a partition of g+s1g+s-1,
@@
$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ $=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{1}}^{\circ}(\pi,(1^{k_{1}}))\times_{\overline{\mathcal{M}}_{1,1}}\ldots\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}})),$
$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ $=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times\overline{\mathcal{M}}^{\circ}_{g_{1}}(\pi,(1^{k_{1}}))\times\ldots\times\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}}))\,$,
where $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ is the base change of $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ along the diagonal $\Delta:\overline{\mathcal{M}}_{1,1}\to\overline{\mathcal{M}}_{1,1}\times\ldots\times\overline{\mathcal{M}}_{1,1}$.
The fixed locus of $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},n,(2,1^{n-2}))$ for the action of the torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ is the disjoint union of the stacks $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$, up to an automorphism factor which we denote by $\operatorname{Aut}(\mathbf{k},\mathbf{g})$, and the restriction of the virtual class $\overline{\mathcal{M}}_{g}^{\bullet}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}$ to $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ agrees with $\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}$. By the localization formula of [19], we have
$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\frac{\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}}{e_{(\mathbb{C}^{*})^{2}}(N^{vir})}\,$,
where the contribution of the virtual normal bundle on each factor of $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ is given by
(3.8) $\begin{split}\frac{1}{\mathsf{e}_{\mathsf{T}}(N^{vir})}&=\frac{c(\mathbb{E}^{\vee}\otimes t_{1})c(\mathbb{E}^{\vee}\otimes t_{2})}{t_{1}t_{2}}\\ &=-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-1}+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-2}-(t_{1}+t_{2})\lambda_{g_{i}-1}\lambda_{g_{i}-2}+\ldots\end{split}$
if $g_{i}>1$, and $\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{1}+1$ if $g_{i}=1$. We denote the class (3.8) by $N_{g_{i}}$, and note that the lower order terms in the Hodge classes will not contribute for dimensional reasons. After integration, we obtain
@@
where $\operatorname{ev}_{i}:\overline{\mathcal{M}}_{(m,\mathbf{k},\mathbf{g})}\to\overline{\mathcal{M}}_{1,1}$ recovers the isomorphism class of the target curve of the $i$-th map. The second equality follows from the decomposition of the diagonal of $\overline{\mathcal{M}}_{1,1}$:
$\Delta]=\sum_{j=0}^{s}[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\boxtimes 1\boxtimes[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\in\operatorname{\mathsf{CH}}^{s}(\overline{\mathcal{M}}_{1,1}\times\cdots\times\overline{\mathcal{M}}_{1,1})\,.$
@@
We now follow the geometric notation and analysis of Section 2.5. The moduli space of stable maps to the fibers of $\pi$ lies over $\overline{\mathcal{M}}_{1,1}$,
@@
[…]
@@
over $\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$. Since $k_{i}>0$, the map (3.10) is injective, so we obtain an exact sequence
@@
where $\mathbb{F}$ is a rank g1g-1 vector bundle on $\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$. We therefore have a factorization
@@
on $\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$, where $E]\in\overline{\mathcal{M}}_{1,1}$ is the moduli point of a fixed nonsingular elliptic curve $E$.
@@
(3.12) $\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})]^{vir}}\lambda_{g_{i}}\lambda_{g_{i}-1}=\frac{1}{24}\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(E,k_{i})]^{vir}}\lambda^{2}_{g_{i}-1}\,.$
@@
Therefore, the only possible non-zero contributions to the right hand side of (3.9) come from triplets $(e,\mathbf{k},\mathbf{g})$ where $\mathbf{g}=(g,1,\ldots,1)$ with $g\geq 1$.
@@ 3.5. Proof of Proposition 8
where the gluing identifies the section $\mathsf{p}_{1}$ of $\mathcal{E}$ with the 0-section of $\mathbb{P}(T_{\mathsf{p}_{1}}\mathcal{E}\oplus\mathbb{C})$. The universal section $\mathsf{p}_{1}$ is now the $\infty$-section of $\pi_{P}$. We will use the degeneration formula
@@ 1. $\bullet$
    $\overline{\mathcal{M}}_{g}(\pi/\mathsf{p}_{1})]^{vir}\cap\operatorname{ev}_{\overline{\mathcal{M}}_{1,1}}^{*}(\mathsf{pt})=[\overline{\mathcal{M}}_{g}(E/p_{1})]^{vir}$
@@ 2. $\bullet$
    Vertex (3.21) vanishes because of the two point conditions over $\overline{\mathcal{M}}_{1,1}$.
@@ 3. $\bullet$
The vertices (3.18) and (3.20) contribution factors 11 and $1/2$ respectively (the latter is cancelled by the multiplicity which occurs in the degeneration formula).
@@ 4.2. Proof of Theorem 4
We adapt the reconstruction strategy of [52, Section 6] to prove the result. Let $m\geq 0$, and fix classes
@@
Pulling back Getzler’s relation from $\overline{\mathcal{M}}_{1,4}$ to $\overline{\mathcal{M}}_{1,4+m}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),d)$ and summing over the curve degree $d$, we obtain the following equation (see [27, Equation (27)])
$3\sum_{h\in S_{4}}\langle v_{h(1)}\star v_{h(2)},v_{h(3)}\star v_{h(4)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\,\sim\,4\sum_{h\in S_{4}}\langle v_{h(1)}\star v_{h(2)}\star v_{h(3)},v_{h(4)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,{\framebox{\tiny{2+m}}}}\,$,
where the relation $\mathsf{X}\sim\mathsf{Y}$ means that the difference $\mathsf{X}-\mathsf{Y}$ is an explicit sum of arbitrary genus 0 invariants and genus $1$ invariants with at most $m+1$ insertions. Getzler’s relation can be equivalently cast in the following form:
@@
from (4.2). A linear algebraic result [52, Lemma 6.2] about solutions to the difference equation then yields
@@
Repeated use of relation (4.7) shows that $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$ can be expressed in terms of genus 0 invariants and $1$-point genus $1$-functions. Moreover, since we use the divisor equation in the argument, the resulting expressions are linear in $1$-point genus $1$ functions and their $q\frac{d}{dq}$ derivatives. By the genus 0 reconstruction result in [39], the genus 0 series can be written as rational functions in matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$. The proof of Theorem 4 is complete. ∎
@@ 3. $\bullet$
Since the eigenvectors $v_{i}$ are found by solving the equation $D\star_{\mathsf{t}}v_{i}=e_{i}v_{i}$, the components of $v_{i}$ are rational functions of $e_{i}$ with coefficients in the matrix coefficients of $D\star_{\mathsf{t}}$. Therefore, the $\Delta_{i}$ are rational functions in the eigenvalues $e_{i}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$, where the matrix coefficients of $D\star_{\mathsf{t}}$ lie.
To evaluate the term $\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})$, we need only evaluate symmetric rational functions in the derivatives of the eigenvalues $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$. By Proposition 14 of Appendix A, these expressions lie in the field of rational functions of derivatives of the symmetric functions of $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$. The outcome is an explicit calculation of
$\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2},q)\,.$
@@
$\nabla_{v_{1}}\nabla_{v_{2}}\cdots\nabla_{v_{k}}\log(\prod_{i}\Delta_{i})\,\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2},q)\,.$